OpenStudy (anonymous):

How would you find the General Solution to the Trigonometric function (Look in replies)

OpenStudy (anonymous):

\[y=arcsin (\frac{- \sqrt3 }{ 2 })\]

OpenStudy (anonymous):

Its _____Plus or Minus 2kPI

OpenStudy (anonymous):

@hartnn

OpenStudy (amistre64):

sin function take an angle and read back a ratio arcsin reverses the process, it takes a ratio and gives back the angle

OpenStudy (amistre64):

so, what arc (angle) has a sine of -sqrt3/2 ?

OpenStudy (anonymous):

uhm. I honestly have no clue. How would I find it?

OpenStudy (amistre64):

you either need to use a calculator that has inverse trig function on it, or if this is one of the special simpler triangle ratios then it would be useful to put those few triangles to memory

OpenStudy (anonymous):

I have a calculator with Inverse trig functions on hand.

OpenStudy (amistre64):



OpenStudy (amistre64):

i recall that an equallateral triangle has the sides we need but the calculator is just as well

OpenStudy (anonymous):

Okay, so how would I input something like this into my calculator to give me the answer I am looking for?

OpenStudy (amistre64):

depends on the calculator really. a ti83 for instance, hit 2nd, sin, and input -sqrt3/2 others ive used would have you input -sqrt3/2 and then press the inverse sine key

OpenStudy (amistre64):

one small issue with the calulator is that its gonna try to calculate the pi part of it into the answer and your not gonna have an 'exact' result;

OpenStudy (amistre64):

the special triangle method, or a unit circle chart are really useful in this case

OpenStudy (anonymous):

I have a unit circle on hand also.

OpenStudy (amistre64):

use the unit circle and find the angle that has a sin value of -sqrt3/2 .... or |sin| = sqrt3/2

OpenStudy (amistre64):



OpenStudy (anonymous):

Would my answer be pi/3? Because the Square root of -3/2 is .866. the Arcsin of .866 is approx. 60.

OpenStudy (amistre64):

almost. pi/3 has a sine of sqrt3/2 -pi/3 has a sine of -sqrt3/2 :)

OpenStudy (anonymous):

My options are pi/3 2pi/3 4pi/3 5pi/3

OpenStudy (amistre64):

well, then we have to work with 2pi instead of 0pi, since 2pi = 0pi in trig stuff 2pi - pi/3 = ?

OpenStudy (anonymous):

aproxx. 5.23

OpenStudy (amistre64):

technically, arcsin is a function that has a range from -pi/2 to pi/2 so -pi/3 is correct and the options are wrong but other than that: 5pi/3 is the same point in the unit circle as -pi/3 so they are equivalent.

OpenStudy (amistre64):

95% of trig stuff is memorization and mental gymnastics ....

OpenStudy (anonymous):

Okay, I see that now. Since 5pi/3 is in quadrant IV, The Sin and Cosin are both negative. Okay okay. I finally understand! Thank you so Much!

OpenStudy (anonymous):

I hate trig. Its a royal pain.

OpenStudy (amistre64):

:) good luck

OpenStudy (anonymous):

can you assist with one more? Its a quick one

OpenStudy (amistre64):

maybe

OpenStudy (anonymous):

Gotta find the Inverse of f(x)=2arcsinx

OpenStudy (anonymous):

The inverse of an inverse :o So wouldnt it be 2sin(x)?

OpenStudy (amistre64):

not quite an inverse function has the property that if f(x) and g(x) are inverses: then f(g) = x f(g(x)) = 2 arcsin(g(x)) arcsin undoes sin so we know there has to be a sin involved 2 arcsin(sin(u)) = 2u when does 2u = x? solve for u

OpenStudy (anonymous):

f-1(x)=sin(x/2)

OpenStudy (amistre64):

correct

OpenStudy (anonymous):

These are my options 1/2 sin(x) sin(1/2x) sin(2x) 2sin(x)

OpenStudy (amistre64):

1/2 x is the same as x/2

OpenStudy (anonymous):

Thanks man :) I appreciate it.

OpenStudy (amistre64):

yw :)